OpenStudy (anonymous):

write an equation of a line that goes through point (4,-5) and is parallel to line y=2x+8.

OpenStudy (cj49):

y − y1 = 2(x − x1) slope y=mx+b here the slope is 2 y-(-5)=2(x-4)

OpenStudy (anonymous):

so y+5=2x-4???

OpenStudy (cj49):

nopes y+5=2x-8

OpenStudy (cj49):

y=2x-8-5 y=2x-13

OpenStudy (anonymous):

oooh, my bad, i forgot to distribute the 2 with the 4.

OpenStudy (anonymous):

so would y=2x-13 be the answer??

OpenStudy (cj49):

yes...

OpenStudy (anonymous):

alright, thank you!! can you help me with one more?

OpenStudy (cj49):

yw..sure

OpenStudy (anonymous):

write an equation of the line in point-slope form, that passes through the given point and has the given slope. point: (-2,15), (9, -18) point: (8, -8); slope 3

OpenStudy (anonymous):

they have those two set of numbers.

OpenStudy (cj49):

so u need 3 equations ryt

OpenStudy (anonymous):

two. one for (-2, 15), 9,-18 and one for (8, -8) slope 3

OpenStudy (cj49):

y-15=-18-15/9-(-2) (x-(-2))

OpenStudy (anonymous):

wpuld i distribute the 9 into all numbers?

OpenStudy (cj49):

y-15=-33/11(x+2) y-15=-3(x+2) y-15=-3x+6 3x+y=21

OpenStudy (anonymous):

ooooh, okay! i see.

OpenStudy (anonymous):

would the one with the slope be y-8=3(x+8)?

OpenStudy (cj49):

y − y1 = 3(x − x1) y-(-8)=3(x-8)

OpenStudy (anonymous):

y+8=3x-8?

OpenStudy (cj49):

yes,combine like terms.

OpenStudy (cj49):

-3x+y=-16

OpenStudy (anonymous):

alright, thank you soooo much!!!

OpenStudy (cj49):

np c: